# GeoVLMath: Enhancing Geometry Reasoning in Vision-Language Models via Cross-Modal Reward for Auxiliary Line Creation

### v1

@@ 5 Related Work
Recent LVLMs (Anthropic, 2025b; DeepMind, 2025a; OpenAI, 2025) have advanced geometric problem solving, especially in plane geometry. The prior methods mainly follow two categories: (i) direct generation of answers or reasoning paths from multimodal inputs (Ning et al., 2025; Xia et al., 2025; Gao et al., 2025), which is limited by intrinsic reasoning capacity; and (ii) tool-augmented reasoning that produces executable code for symbolic computation (Zhao et al., 2025; Sharma et al., 2025; Chen et al., 2024), offloading difficult steps to external engines. However, most methods still handle visual and textual modalities largely separately and lack mechanisms to incorporate auxiliary lines, which restricts performance on tasks requiring such constructions. Visual Sketchpad (Hu et al., 2024) prompts LVLMs to generate code for auxiliary lines that are rendered into augmented diagrams, but relies on precise code and exact coordinate annotations that are rare in real diagrams. In contrast, our approach integrates auxiliary-line construction into a RL framework guided by a cross-modal reward model, decouples construction from the LVLM’s intrinsic reasoning, generalizes across model scales, and removes the need for explicit coordinates, enabling robust auxiliary-line reasoning in solid geometry. Details on benchmarks and datasets are provided in Appendix F.
@@ 6 Conclusion
We present a framework for solving solid geometry problems that require auxiliary lines, an underexplored area in vision-language reasoning. To support this task, we curate AuxSolidMath, a high-quality dataset from real high-school exams with paired diagrams, problem statements, auxiliary-line descriptions, and final answers, enabling both answer supervision and vision-language reward modeling. At the core is a vision-based reward model that scores the agreement between generated auxiliary-line descriptions and ground-truth auxiliary-line-annotated diagrams, providing stable signals for reinforcement learning. Using this dataset and reward model, we train GeoVLMath, an open-source LVLM optimized with RL. Experiments demonstrate state-of-the-art results among open-source LVLMs of comparable size and competitive performance against advanced closed-source LVLMs such as GPT-4o. We will make AuxSolidMath and GeoVLMath publicly available on Hugging Face to facilitate reproducibility, benchmarking, and further research.
Appendix A The Use of Large Language Models
In this paper, the authors used ChatGPT solely for language polishing, including grammar, phrasing, and stylistic refinement. We did not use it to generate scientific content, such as research ideas, methods, or related work. We did not provide any confidential, personal, or proprietary data to the model. The authors take full responsibility for all scientific content, which was exclusively written and verified by us.
Appendix B Limitation and Future Work
Although our framework achieves competitive performance, it does not yet realize the ideal approach of directly rendering precise auxiliary lines on the diagram. Given the limited geometric controllability of current image editing and diffusion models, we instead employ a cross-modal reward model as a practical proxy to promote diagram-text consistency without directly editing the diagram images. In future work, we will investigate constraint-guided diffusion in conjunction with a geometry-constrained rendering engine to explicitly render auxiliary lines, thereby enhancing the alignment between visual constructions and symbolic reasoning.
Appendix C Qualitative Comparison of Image Editing Models for Auxiliary-Line Generation
To further highlight the limitations of current image-editing models in precise geometric construction, we present a single illustrative example comparing two representative models on a three-dimensional coordinate system construction task. As shown in Figure 5, this comparative example reveals a persistent difficulty in faithfully instantiating the specified auxiliary line descriptions, namely, aligning the edits with the intended spatial constraints, which motivates our text-driven auxiliary line construction guided by a cross-modal reward model.
Figure 5: Comparison of two representative image editing models for constructing a three-dimensional Cartesian coordinate system.
Appendix D Data Creation
@@
D.1 Automated Problem Identification
@@
D.2 Automated Deduplication and Diagram Extraction
@@
D.3 Structured Data Extraction
Building on the high-quality geometric diagrams obtained in the previous step, we proceed to extract the corresponding textual content for each geometry instance, including the problem description, the final answer, and the auxiliary-line description. This extraction process is non-trivial, as the original Word documents frequently embed mathematical expressions using MathType formats that are not reliably handled by standard document parsing tools.
To address this challenge, we render the processed Word documents as images, thereby enabling LVLMs to leverage their visual reasoning capabilities. Although this approach appears straightforward, open-source models such as Qwen2.5-VL-7B-Instruct (Bai et al., 2025) often struggle to accurately parse complex geometry problems involving symbolic notation and mathematical expressions. In contrast, closed-source models like GPT-4o (Hurst et al., 2024) exhibit significantly stronger performance, but their reliance on commercial APIs introduces substantial costs and limits scalability in large-scale applications. To balance accuracy with scalability, we adopt a hybrid strategy. More concretely, we first utilize an advanced closed-source model (i.e., GPT-4o) to generate a small, high-quality seed dataset comprising 300 manually verified instances. This curated dataset is then used to fine-tune an open-source LVLM (i.e., Qwen2.5-VL-7B-Instruct), resulting in a lightweight, domain-adapted model capable of accurate and scalable text extraction. The final output consists of the extracted problem description, the final answer, and the auxiliary-line description, all encapsulated in a structured JSON format. This unified representation facilitates consistent data handling and serves as a foundation for training a robust open-source text extraction model. By releasing this model, we aim to contribute a practical and reusable resource to the broader research community working on geometry-aware vision-language understanding.
D.4 Manual Verification
@@
Appendix E Experimental Setup
E.1 Models
On the closed-source models, we include leading models such as gpt-5-mini (OpenAI, 2025), o4-mini (OpenAI, 2025) and GPT-4o (Hurst et al., 2024), Gemini-2.0-Flash and Gemini-2.5-Flash (DeepMind, 2025a), Claude 3.7 Sonnet (Anthropic, 2025a) and Claude Sonnet 4 20250514 (Anthropic, 2025b). These models represent the forefront of multimodal reasoning among closed-source models, although their internal architectures remain undisclosed. On the open-source models, we consider several publicly available high-performance models, including the Qwen2 VL (Wang et al., 2024b) and Qwen2.5 VL series (Bai et al., 2025), InternVL 3 families (Zhu et al., 2025), LLaMA-3.2-11B-Vision-Instruct (Meta, 2024) and Llama-4-Scout-17B-16E-Instruct (Meta, 2025). These models encompass a range of design paradigms, parameter scales, and instruction tuning strategies, providing a robust comparative foundation for evaluating multimodal reasoning capabilities. Note that models such as Gemini-2.5 Pro (DeepMind, 2025b) and OpenAI o3 (OpenAI, 2025) are excluded from our study due to limited accessibility and high inference costs.
E.2 Training Implementation Details
@@
RL Stage. The RL phase is performed using the EasyR1 framework (Zheng et al., 2025) with the GRPO algorithm. For the Qwen2.5-VL-7B model, training and validation data are loaded from Parquet files containing question-diagram pairs, with a maximum response length of 8192. Both rollout and validation batch size are set to 16. The actor is optimized using AdamW (learning rate 2e-6, weight decay 1e-2, no warmup). KL regularization is applied using the low_var_kl penalty with a coefficient of 1e-2. Training runs for 6 epochs using bf16 precision, with gradient checkpointing and partial FSDP offloading enabled for memory efficiency.
@@
All training was conducted on a server equipped with two NVIDIA A100 80GB and two NVIDIA A800 80GB GPUs. The SFT stage was performed on the A100 GPUs, while the full set of four GPUs was utilized during the reinforcement learning stage.
Appendix F Related Work
@@
In this section, we also review benchmarks and datasets for geometric reasoning. Most benchmarks and datasets for geometric problem solving focus on plane geometry, where diagrams and problems involve two-dimensional figures. Well-known resources in this area include Geometry3K (Lu et al., 2021), GeoQA (Chen et al., 2021a), UniGeo (Chen et al., 2022), and GeomRel (Wang et al., 2025b), which primarily cover plane geometry problems. A concurrent benchmark, GeoLaux (Fu et al., 2025), explores the use of auxiliary lines in plane geometry, but is limited to simple cases and lacks engagement with the spatial complexity of solid geometry. Nonetheless, there remains a lack of dedicated resources for solid geometry, even though solving such problems often requires interpreting three-dimensional relationships and drawing auxiliary lines to uncover hidden spatial structures. While SolidGeo (Wang et al., 2025a) is a recent benchmark that focuses exclusively on solid geometry, it does not explicitly require auxiliary lines for solving its problems, leaving this important aspect of spatial reasoning underexplored. Similarly, other benchmarks such as MathVista (Lu et al., 2024), MathVision(Wang et al., 2024a), and MathVerse (Zhang et al., 2024) contain only a limited number of solid geometry problems, and these also do not require auxiliary lines to reach the solution. As a result, these resources fall short of evaluating a model’s ability to solve complex solid geometry problems where auxiliary lines are essential for uncovering implicit spatial relationships. To address this gap, we present AuxSolidMath, the first dedicated dataset for solid geometry problems that require auxiliary lines to solve. It offers comprehensive multimodal supervision, including the original diagram, the problem statement, textual descriptions of the required auxiliary lines, the final answer, and a corresponding diagram annotated with those lines, enabling models to learn how the auxiliary lines facilitate solid geometry reasoning.
Appendix G Prompts
G.1 Prompts for Supervised Fine-tuning
The following presents the two-part prompt template used in our supervised dataset. The system prompt assigns the solver role and enforces formatting: auxiliary lines must be wrapped in [AUX]…[/AUX] and the final answer must appear as plain text in Final Answer:.... The user prompt is multimodal, pairing a diagram referenced by the <image> token with the natural language question {question}, which yields explicit reasoning steps and a final answer.
@@ G.1 Prompts for Supervised Fine-tuning
SYSTEM_PROMPT_FOR_SFT SYSTEM_PROMPT_FOR_SFT = """ You are a mathematician skilled in solving geometry problems through step-by-step reasoning. Solve the given geometry problem based on a geometric diagram and a natural language question. Use ‘[AUX]…[/AUX]’ to indicate auxiliary constructions, such as establishing coordinate systems or constructing auxiliary lines. Finally, provide your final answer within ‘Final Answer:…’. """
USER_PROMPT_FOR_SFT USER_PROMPT_FOR_SFT = """ Image: <image> Question: {question} """
G.2 Prompts for Cross-Modal Reward Model
Using the prompt below, the cross-modal reward model compares the description of the auxiliary line generated by the policy model against a pair of diagrams, the original image $I$ and its auxiliary-line counterpart $I^{+}$, and returns a single line justification and a calibrated score in $[0,1]$ that measures visual–textual agreement. The instruction emphasizes the correctness of auxiliary-line constructions and adherence to geometric constraints. Higher scores indicate stronger alignment.
@@ G.2 Prompts for Cross-Modal Reward Model
Appendix H Representative Examples
@@
Appendix I Case Study
I.1 Successful Cases

### v2

@@ 5 Related Work
Recent LVLMs (Anthropic, 2025b; DeepMind, 2025b; OpenAI, 2025b) have advanced geometric problem solving, particularly in plane geometry. Prior work mainly follows two lines: direct generation of answers or reasoning paths from multimodal inputs (Ning et al., 2025; Xia et al., 2025; Gao et al., 2025), which is constrained by intrinsic reasoning capacity, and tool-augmented reasoning that generates executable code for symbolic computation or geometric operations (Zhao et al., 2025; Sharma et al., 2025; Chen et al., 2024). More recently, a code-driven visual construction paradigm has emerged, where models generate code to construct or modify diagrams during reasoning, as exemplified by Visual Sketchpad (Hu et al., 2024), V-Thinker (Qiao et al., 2025), and CodePlot-CoT (Duan et al., 2025). While these methods enable interactive visual feedback, they depend on precise code execution and explicit coordinate annotations, and tightly couple auxiliary-line construction with the LVLM’s intrinsic reasoning, thereby limiting robustness in solid geometry. In contrast, our approach incorporates auxiliary-line construction into an RL framework guided by a cross-modal reward model, decoupling construction quality from intrinsic reasoning, without explicit geometric specifications, enabling robust auxiliary-line reasoning across model scales.
@@ 6 Conclusion
Auxiliary-line reasoning in solid geometry exposes a key limitation of current LVLMs: the lack of reliable learning signals for structured visual construction. We address this challenge by formulating auxiliary-line construction as a vision-language alignment problem and introducing a RL framework guided by a cross-modal reward that directly measures diagram-text correspondence. This design enables stable optimization without relying on explicit geometric specifications or executable code. To support learning at scale, we automatically construct AuxSolidMath, a high-quality dataset of real-exam solid geometry problems with aligned diagrams and auxiliary-line annotations, providing a reusable resource for the community. Leveraging this framework, GeoVLMath consistently improves auxiliary-line reasoning.
Limitations
While our approach demonstrates strong effectiveness for auxiliary-line reasoning in solid geometry, our experiments primarily consider settings where auxiliary-line constructions admit concise high-level descriptions. Extending the evaluation to even broader or more diverse geometric constructions remains an interesting direction for future work. In addition, our cross-modal reward model relies on annotated auxiliary-line diagrams for supervision during reinforcement learning. Although we introduce AuxSolidMath to support this setting, exploring more scalable forms of weak or self-supervised reward signals is a promising direction for future research.
Appendix A The Use of AI Assistants
In this paper, ChatGPT was used exclusively for language polishing, including grammar correction, phrasing, and stylistic refinement. It was not used to generate scientific content such as research ideas, methods, experiments, or related work. No confidential, personal, or proprietary information was shared with the model. The authors take full responsibility for the scientific content, which was entirely authored and verified by the authors.
Appendix B Data Creation
@@
B.1 Automated Problem Identification
@@
B.2 Automated Deduplication and Diagram Extraction
@@
B.3 Structured Data Extraction
Building on the high-quality geometric diagrams obtained in the previous step, we proceed to extract the corresponding textual content for each instance, including the problem description, the final answer, and the auxiliary-line description. This extraction process is non-trivial, as the original Word documents frequently embed mathematical expressions in MathType formats that are not reliably supported by standard document parsing tools.
To address this challenge, we render the processed Word documents as images, thereby enabling LVLMs to leverage their visual reasoning capabilities. Although this approach appears straightforward, open-source models such as Qwen2.5-VL-7B-Instruct (Bai et al., 2025) often struggle to accurately parse complex geometry problems involving symbolic notation and mathematical expressions. In contrast, closed-source models like GPT-4o (Hurst et al., 2024) exhibit significantly stronger performance, but their reliance on commercial APIs introduces substantial costs and limits scalability in large-scale applications. To balance accuracy with scalability, we adopt a hybrid strategy [^4] DeepSeek-OCRWei et al. (2025) was released on October 20, 2025 and was not available during data construction. A Qwen2.5-VL-based extractor fine-tuned on 300 seed samples achieved approximately 98% accuracy and was therefore used.. Specifically, we first utilize an advanced closed-source model (i.e., GPT-4o) to generate a small, high-quality seed dataset comprising 300 manually verified instances. This curated dataset is then used to fine-tune an LVLM (i.e., Qwen2.5-VL-7B-Instruct), resulting in a lightweight, domain-adapted model capable of accurate and scalable text extraction. The final output consists of the extracted problem description, the final answer, and the auxiliary-line description, all encapsulated in a structured JSON format. This unified representation facilitates consistent data handling and serves as a foundation for training a robust open-source text extraction model. By releasing this model, we aim to contribute a practical and reusable resource to the broader research community working on geometry-aware vision-language understanding.
B.4 Manual Verification
@@
Appendix C Experimental Setup
C.1 Models
On the closed-source models, we include leading models such as gpt-5-mini (OpenAI, 2025a), o4-mini (OpenAI, 2025b) and GPT-4o (Hurst et al., 2024), Gemini-2.0-Flash and Gemini-2.5-Flash (DeepMind, 2025b), Claude 3.7 Sonnet (Anthropic, 2025a) and Claude Sonnet 4 20250514 (Anthropic, 2025b). These models represent the forefront of multimodal reasoning among closed-source models, although their internal architectures remain undisclosed. On the open-source models, we consider several publicly available high-performance models, including the Qwen2 VL (Wang et al., 2024b) and Qwen2.5 VL series (Bai et al., 2025), InternVL 3 families (Zhu et al., 2025), LLaMA-3.2-11B-Vision-Instruct (Meta, 2024) and Llama-4-Scout-17B-16E-Instruct (Meta, 2025). In addition, we include representative methods following the code-driven visual construction paradigm, namely V-Thinker (Qiao et al., 2025) and CodePlot-CoT (Duan et al., 2025), which generate executable code to construct or modify diagrams as part of the reasoning process. These approaches provide an important comparison point for evaluating different strategies of integrating visual construction into multimodal reasoning. Together, these models encompass a range of design paradigms, parameter scales, and instruction tuning strategies, forming a comprehensive and robust foundation for evaluating auxiliary-line reasoning in multimodal settings. Note that models such as Gemini-2.5 Pro (DeepMind, 2025c) and OpenAI o3 (OpenAI, 2025b) are excluded from our study due to limited accessibility and high inference costs.
C.2 Training Implementation Details
@@
RL Stage. For the Qwen2.5-VL-7B model, both training and validation data are loaded from Parquet files containing question-diagram pairs. The maximum response length is set to 8,192 tokens, and both the rollout and validation batch sizes are set to 16. The actor is optimized using AdamW (learning rate 2e-6, weight decay 1e-2, no warmup). KL regularization is applied using the low_var_kl penalty with a coefficient of 1e-2. Training runs for 6 epochs using bf16 precision, with gradient checkpointing and partial FSDP offloading enabled for memory efficiency.
@@
All training was conducted on a server equipped with two NVIDIA A100 80GB and two NVIDIA A800 80GB GPUs. The SFT stage was performed on two A100 GPUs, whereas the reinforcement learning stage utilized all four GPUs. For response generation during training and evaluation, we enabled stochastic sampling with temperature set to 0.7 and top-p set to 0.95.
Appendix D Related Work
@@
In this section, we also review benchmarks and datasets for geometric reasoning. Most benchmarks and datasets for geometric problem solving focus on plane geometry, where diagrams and problems involve two-dimensional figures. Well-known resources in this area include Geometry3K (Lu et al., 2021), GeoQA (Chen et al., 2021a), UniGeo (Chen et al., 2022), and GeomRel (Wang et al., 2025b), which primarily cover plane geometry problems. A concurrent benchmark, GeoLaux (Fu et al., 2025), explores the use of auxiliary lines in plane geometry, but is limited to simple cases and lacks engagement with the spatial complexity of solid geometry. Nonetheless, there remains a lack of dedicated resources for solid geometry, even though solving such problems often requires interpreting three-dimensional relationships and drawing auxiliary lines to uncover hidden spatial structures. While SolidGeo (Wang et al., 2025a) is a recent benchmark that focuses exclusively on solid geometry, it does not explicitly require auxiliary lines for solving its problems, leaving this important aspect of spatial reasoning underexplored. Similarly, other benchmarks such as MathVision(Wang et al., 2024a), MathVista (Lu et al., 2024), and MathVerse (Zhang et al., 2024) contain only a limited number of solid geometry problems, and these also do not require auxiliary lines to reach the solution. As a result, these resources fall short of evaluating a model’s ability to solve complex solid geometry problems where auxiliary lines are essential for uncovering implicit spatial relationships. To address this gap, we present AuxSolidMath, the first dedicated dataset for solid geometry problems that require auxiliary lines to solve. It offers comprehensive multimodal supervision, including the original diagram, the problem statement, textual descriptions of the required auxiliary lines, the final answer, and a corresponding diagram annotated with those lines, enabling models to learn how the auxiliary lines facilitate solid geometry reasoning.
@@
SYSTEM_PROMPT_FOR_SUPERVISED_FINE-TUNING SYSTEM_PROMPT_FOR_SUPERVISED_FINE-TUNING = """ You are a mathematician skilled in solving geometry problems through step-by-step reasoning. Solve the given geometry problem based on a geometric diagram and a natural language question. Use ‘[AUX]…[/AUX]’ to indicate auxiliary constructions, such as establishing coordinate systems or constructing auxiliary lines. Finally, provide your final answer within ‘Final Answer:…’. """
USER_PROMPT_FOR_SUPERVISED_FINE-TUNING USER_PROMPT_FOR_SUPERVISED_FINE-TUNING = """ Image: <image> Question: {question} """
Figure 5: Prompts for supervised fine-tuning.
@@
Figure 6: Prompts for cross-modal reward model.
Appendix E Prompts
E.1 Prompts for Supervised Fine-tuning
Figure 5 presents the two-part prompt template used in our supervised dataset. The system prompt assigns the solver role and enforces formatting: auxiliary lines must be wrapped in [AUX]…[/AUX] and the final answer must appear as plain text in Final Answer:.... The user prompt is multimodal, pairing a diagram referenced by the <image> token with the natural language question {question}, which yields explicit reasoning steps and a final answer.
E.2 Prompts for Cross-Modal Reward Model
Using the prompt as shown in Figure 6, the cross-modal reward model compares the description of the auxiliary line generated by the policy model against a pair of diagrams, the original image $I$ and its auxiliary-line counterpart $I^{+}$, and returns a single line justification and a calibrated score in $[0,1]$ that measures visual–textual agreement. The instruction emphasizes the correctness of auxiliary-line constructions and adherence to geometric constraints. Higher scores indicate stronger alignment.
Appendix F Representative Examples
@@
Appendix G Case Study
G.1 Successful Cases